OpenStudy (anonymous):

If (x + 1)(x - 3) = 5, then which of the following statements is true? x + 1 = 0 or x - 3 = 0 x + 1 = 5 or x - 3 = 5 x - 4 = 0 or x + 2 = 0

OpenStudy (anonymous):

expanding you get x^2-2x-3=5 so x^2-2x-8=0 so (x-4)(x+2)=0 then x-4=0 or x+2=0

OpenStudy (anonymous):

ok thanks you

OpenStudy (anonymous):

no problem